User name Elena Kh.
Log entry time 08:07:46 on May17,2007
Entry number 202425
People on shift:
Elena Khrosinkova - SL,
Michael Olson - TO
00:52 - Michael found that we lost the connection to iocha4, could not connect to hareboot2, crate resets for hadron detector stack were blank, called Brad Sawatzky -> reboot Right Arm Detector' Slow Controls VME Reset: iocha4, hadron crates reset came back. Later rebooted all VDC high voltages.
04:37 - 05:00 - beam off, problems with hallC beam.
05:41 - 05:52 - changed the target to Dummy, 20 cm
3537 - 3551 - 45bQ2, He3
3552 - 3554 - 45bQ2, Dummy, 20 cm
The runs with some problems but not junk:
3537 - 3542 - without connection to iocha4, evrything else looks fine.
3540 - 3544 - no reading for some of the VDC HV (Start of Run ...)